THE LABORATORY DETECTION AND GEOMETRICAL STRUCTURE OF THIOZONE (S_{3}) AND TETRASULFUR (S_{4})

\underline{M. C. McCARTHY}, S. THORWIRTH, C. A. GOTTLIEB, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A.

Small sulfur clusters (S_{2} – S_{8}) are of significant fundamental and applied interest, but surprisingly little is known about the electronic ground state and structures from S_{3} to S_{5}. The rotational spectra of two, thiozone, S_{3}, and tetrasulfur, S_{4}, have now been observed for the first time by Fourier transform microwave spectroscopy. From the rotational constants of the normal and ^{34}S isotopic species, precise geometrical structures have been derived for both clusters with C_{2v} symmetry. The derived structural parameters indicate substantial double bonding in S_{3} and sp^{2} hybridization of the apex sulfur atom; S_{4} also has double bonding between the outermost S atoms, but the central bond is relatively long and weak, like that observed for S_{2}O_{2}.